OpenStudy (anonymous):

The vapor pressure of ethanol at 25 degrees C is 7.83kPa . Calculate the vapor pressure in atm. Round answer to 3 significant digits.

OpenStudy (anonymous):

1 Pascal equals 9.8692 * 10^-6 atmospheres 7.83 kPa = 7.83 * 1000 Pa = 78300 Pa

OpenStudy (anonymous):

........

OpenStudy (anonymous):

So is it a = amount, t = tempurature, m = measure? A is 78300, t is 25 and m is what we are finding?

OpenStudy (anonymous):

Mymathhomework this is not a question on exponential decay formula, or, for that matter, any formula. It is only a unit conversion question

OpenStudy (anonymous):

So 78300 is the answer? What is the atm for?

OpenStudy (anonymous):

Pa means 'pascals' and atm means 'atmospheres', both are units of pressure

OpenStudy (anonymous):

So I don't need to use the 25 degrees C any where?

OpenStudy (anonymous):

As far I I know I don't think the 25 degree C is relevant to the conversion

OpenStudy (anonymous):

OK, so it says to round to 3 significant digits so I will need to put it in scientific notation like this: 7.83 X 10^4 right? or is the 4 negative?

OpenStudy (anonymous):

First you get 7.83 kPa = 7.83 * 10^4 Pa Then convert it to atm by using the relation 1 Pa = 9.8692 * 10^-6 atm, this gives you 7.83 kPa = 7.83 * 10^4 * 9.8692 * 10^-6 atm = 77.275836 * 10^-2 atm = 0.77275836 atm Now round to three significant digits, = 0.773 atm